Water Use Calculation Problem:

In the following problem calculate the water savings for Shower Use using two different GPM shower heads and then calculate the water savings for Toilet Use using two different GPFlush toilets. The last answer in the problem is to calculate the total combined water savings realized by both Shower Use Savings and Toilet Flush Savings together

C. Calculate the total gallons of water savings for Shower Use Water by reducing the GPM

Total Gallons Saved ________________
Answer #1

A) Savings of water using shower heads with two different flow rates:

Let's take two shower heads with flow rates 2.5 GPM and 1.5 GPM

And consider diaily water running time is 1hr, then

  • Water usage with 2.5 GPM shower head = ( water flow rate of shower head) * ( Running time) = 2.5 * 60 = 150 gallon
  • Water usage with 1.5 GPM shower head = ( water flow rate of shower head) * ( Running time) = 1.5 * 60 = 90 gallon
  • Therefore daily water saving using second shower head over first is = 150 - 90 = 60 gallon.

B) Savings of water using two different GPFlush toilets:

Let's take two toilets with 1.5 GPFiush and 1 GPFlush

Consider total average daily flushes are 10, then

  • Water usge by 1.5 GPFlush toilet = 1.5 (gallon) * 10 (flushes) = 15 gallon
  • Water usage by 1 GPFlush toilet = 1 (gallon) * 10 (flushes) = 10 gallon
  • Therefore daily savings of water using second toilet over first = 15 - 10 = 5 gallon.

Last answer : Combined water savings

realized by both Shower Use Savings and Toilet Flush Savings together = 60 + 5 = 65 gallon.

C)Daily total gallons of water savings for Shower Use Water by reducing the GPM = 60 gallon